# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v1

@@ 4 The segment neighbourhood property (SNP) and some metric lemmas
By the previous lemma, a projection of $b_{1}(0)$ on $b_{2}$ lies in $[b_{2}(3R),b_{2}(\mathrm{L}(b_{2})-3R]$. If such projection lies in $[b_{2}(R),b_{2}(3R)]$ or $[b_{2}(\mathrm{L}(b_{2})-3R),b_{2}(\mathrm{L}(b_{2})-R)]$, then $\gamma$ has a point which is $5R$-close to either $b_{2}(3R)$ or $b_{2}(\mathrm{L}(b_{2})-3R)$ by the (SNP), otherwise the same conclusion follows by Lemma 4.6. ∎
@@
Let $(X,\mathsf{d})$ be a compact metric space satisfying the (SNP) (see Definition 4.2). Let $l$ be a segment of maximal length in $X$ and assume that $\mathrm{L}(l)\geq 200R$. If $X$ is not contained in a $200R$-neighbourhood of $l$, then there exists $L>0$ and a distance preserving map $\phi:S^{1}_{L}\to X$, such that $X$ is contained in the $D$-neighbourhood of $\phi(S^{1}_{L})$. In particular, $L\geq 50R$.
@@
Consider now a sequence $\phi_{i}\in A$ minimizing the length functional. Each image of $\phi_{i}$ has a projection on $\gamma_{1}$ lying in $I_{1}^{\gamma_{1}}$, so that by the (SNP), it has a point which is $D$-close to $I_{1}^{\gamma_{1}}$. Hence, all the images of the maps $\phi_{i}$ are contained in a bounded set of $X$. Since the $\phi_{i}$ are all $1$-Lipschitz, by Ascoli-Arzelà theorem, there exists (modulo passing to a subsequence) $\phi_{\infty}$ such that $\phi_{i}\to\phi_{\infty}$ uniformly. By our previous point $\phi_{\infty}\in A$. Since the length functional is lower semicontinuous w.r.t. uniform convergence, $\phi_{\infty}$ is an element of minimal length in $A$.
@@
Let $b_{1},b_{2}\in\{\gamma_{1},\gamma_{2},l\}$ with $b_{1}\neq b_{2}$. Let $t^{b_{1}},t^{b_{2}}\in[0,T]$ be the numbers given in the definition of $A$ relative to $\phi_{\infty}$. Modulo reparametrizing $\phi_{\infty}$ and $\gamma_{1}+\gamma_{2}+l$, we can assume that $b_{1}(\mathrm{L}(b_{1}))=b_{2}(0)$ and $t^{b_{1}}\leq t^{b_{2}}$ with $t^{b_{3}}\notin[t^{b_{1}},t^{b_{2}}]$, i.e. $[t^{b_{1}},t^{b_{2}}]$ corresponds to the arc of $S^{1}_{L}$ that connects $t^{b_{1}}$ and $t^{b_{2}}$ without crossing $t_{b_{3}}$.
@@
By definition of $\phi_{\infty}$, the restriction ${\phi_{\infty}}_{|[t^{b_{1}},t^{b_{2}}]}$ is a segment. By Lemmas 4.14 and 4.6 (see also Remark 4.4), $\pi_{b_{1}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of either $[b_{1}(3R),b_{1}(\bar{t}^{b_{1}})]$ or $[b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$. If the former case happens, ${\phi_{\infty}}([t^{b_{1}},t^{b_{2}}])$ has a point which is $7D$-close to $b_{1}(3R)$, so that it also contains a point $q$ which is $5R$-close to $b_{1}(0)$. By Corollary 4.15, $b_{1}(0)$ is at least $\min\{\mathrm{L}(b_{1}),\mathrm{L}(b_{3})\}+\mathrm{L}(b_{2})/2-20R$ distant from $\phi_{\infty}(t^{b_{2}})$. Hence,
@@
Modulo reparametrizing, we assume $b_{1}(\mathrm{L}(b_{1}))=b_{2}(0)$ and $t^{b_{1}}\leq t^{b_{2}}\leq t^{b_{3}}$ in $[0,T]$. By Step $1$, there exist $t_{1,2}\in[t^{b_{1}},t^{b_{2}}]$ and $t_{2,3}\in[t^{b_{2}},t^{b_{3}}]$, such that $\phi_{\infty}(t_{1,2})$ is $7D$-close to $b_{2}(\mathrm{L}(b_{2})-20R)$ and $\phi_{\infty}(t_{2,3})$ is $7D$-close to $b_{2}(\mathrm{L}(b_{2})+20R)$. Let $\gamma$ be a segment connecting $\phi_{\infty}(t_{1,2})$ and $\phi_{\infty}(t_{2,3})$. By item 2 of Lemma 4.6 and the (SNP), either $\pi_{b_{2}}(\gamma)$ intersects $I^{b_{2}}$, or $\gamma$ contains points that are $7D$ close to $b_{2}(R)$ and $b_{2}(\mathrm{L}(b_{2})-R)$. In the latter case, since $\mathrm{L}(b_{2})\geq 200R$, $\gamma$ would not be a segment. Hence, $\pi_{b_{2}}(\gamma)$ intersects $I^{b_{2}}$. It follows that $\gamma+{\phi_{\infty}}_{|S^{1}_{L}\setminus[t_{1,2},t_{2,3}]}$ belongs to $A$. In particular, ${\phi_{\infty}}_{|[t_{1,2},t_{2,3}]}$ is shorter than $\gamma$, so that it is itself a segment.
@@
It follows that $\phi_{\infty}:S^{1}_{L}\to X$ preserves distances. Since $2\pi L>2R$, the (SNP) implies that $X$ is in the $D$-neighbourhood of $\phi_{\infty}(S^{1}_{L})$. Since $X$ contains a segment of length greater than $200R$, it follows that $L\geq 50R$, concluding the proof. ∎
The next theorem, summarizes the results of this section.
@@
Let $(X,\mathsf{d})$ be a metric space satisfying the (SNP). There exists a $1$-dimensional manifold $I$, possibly with boundary, and a distance preserving map $\phi:I\to X$, such that $X$ is contained in the $200R$-neighbourhood of $\phi(I)$.
@@ 5 Large scale geometry of almost non-negative Ricci and integrally-positive scalar curvature
5.1 Sufficient curvature conditions for (SNP)
@@
Let $r$ be a segment of length greater than $2R$ in $X$. For every $t\in(R,\mathrm{length}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$, i.e., $(X,\mathsf{d})$ satisfies the segment neighbourhood property (SNP) (see Definition 4.2).
In particular, there exists a $1$-dimensional connected manifold $I$ (possibly with boundary) and a distance preserving map $\phi:I\to X$ such that $X$ is contained in the $C(D,R)$-neighbourhood of $\phi(I)$.
@@
Finally, by Theorem 4.17 , there exists a $1$-dimensional connected manifold $I$ (possibly with boundary) and a distance preserving map $\phi:I\to X$ such that $X$ is contained in the $C(D,R)$-neighbourhood of $\phi(I)$. ∎
@@
Let $r$ be a segment of length greater than $2R$ in $X$. For every $t\in(R,\mathrm{length}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$, i.e., $(X,\mathsf{d})$ satisfies the segment neighbourhood property (SNP) (see Definition 4.2).
In particular, there exists a $1$-dimensional connected manifold $I$ (possibly with boundary) and a distance preserving map $\phi:I\to X$ such that $X$ is contained in the $C(D,R)$-neighbourhood of $\phi(I)$.
The remaining part of the section is devoted to showing that if a manifold satisfies one of the conditions (34), (35), (36), so do its coverings (modulo a rescaling by dimensional constants). This is the content of Corollaries 5.5.
@@
Let $n\in\mathbb{N}$ and let $K\in\mathbb{R}$. Let $(M^{n},g,p)$ be a pointed Riemannian manifold with $\mathsf{Ric}_{M}\geq K$. Let $(\tilde{M}^{n},\tilde{g},\tilde{p})$ be a Riemannian covering, with covering map $\pi:\tilde{M}\to M$ such that $\pi(\tilde{p})=p$. Let $F:M\to\mathbb{R}_{+}$ be a measurable locally bounded function and let $\tilde{F}:=F\circ\pi$.
@@
    $\fint_{\tilde{B}_{5}(\tilde{p})}\tilde{F}\,{\rm d}\widetilde{\mathsf{Vol}}\geq c(K,n)\fint_{B_{1}(p)}F\,{\rm d}\mathsf{Vol}.$
@@
    $\fint_{\tilde{B}_{5}(\tilde{p})}\tilde{F}\,{\rm d}\widetilde{\mathsf{Vol}}\leq c(K,n)\fint_{B_{1}(p^{\prime})}F\,{\rm d}\mathsf{Vol}.$
@@
We first claim that
$\max_{x\in B_{1}(p)}\#(\pi^{-1}(x)\cap\tilde{B}_{1}(\tilde{p}))\leq\min_{y\in B_{1}(p)}\#(\pi^{-1}(y)\cap\tilde{B}_{3}(\tilde{p}))<+\infty.$ (41)
To prove the claim, let $x,y\in B_{1}(p)$. Since $\pi$ is a covering map, #$\#(\pi^{-1}(x)\cap\tilde{B}_{1}(\tilde{p}))$ is finite and positive. Let $l\in\mathbb{N}$ be such that
$\pi^{-1}(x)\cap\tilde{B}_{1}(\tilde{p})=\{\tilde{x}_{1},\cdots,\tilde{x}_{l}\}.$
We recall that $M$ is the quotient of $\tilde{M}$ by a group of isometries acting in a proper and discontinuous way, and that $\pi:\tilde{M}\to M$ is the projection on such a quotient. Let $\gamma_{1},\cdots,\gamma_{l}$ be the isometries of $\tilde{M}$ (as above) such that $\gamma_{i}(\tilde{x}_{i})=\tilde{x}_{1}$. Let $\tilde{y}\in\pi^{-1}(y)\cap\tilde{B}_{1}(\tilde{p})$ ($\tilde{y}$ exists, as before, because $\pi$ is a covering map). Denoting by $\tilde{\mathsf{d}}$ the distance induced by $\tilde{g}$ on $\tilde{M}$, then
$\tilde{\mathsf{d}}(\gamma_{i}(\tilde{y}),\tilde{p})\leq\tilde{\mathsf{d}}(\gamma_{i}(\tilde{y}),\gamma_{i}(\tilde{x}_{i}))+\tilde{\mathsf{d}}(\gamma_{i}(\tilde{x}_{i}),\tilde{p})=\tilde{\mathsf{d}}(\tilde{y},\tilde{x}_{i})+\tilde{\mathsf{d}}(\tilde{x}_{1},\tilde{p})\leq 3.$
Hence, $\{\gamma_{1}(\tilde{y}),\cdots,\gamma_{l}(\tilde{y})\}\subset\pi^{-1}(y)\cap\tilde{B}_{3}(\tilde{p})$. Since $\gamma_{i}(\tilde{y})\neq\gamma_{j}(\tilde{y})$ if $i\neq j$, the claim (41) follows.
Proof of 1. Let $\{A_{1},\cdots,A_{l}\}$ be a partition of $B_{1}(p)$ such that each $A_{i}$ is contained in a covering chart of $\pi$. For every $i\in\{1,\cdots,l\}$, we set
$N_{i}:=\max_{y\in A_{i}}\,\#(\pi^{-1}(y)\cap\tilde{B}_{3}(\tilde{p})).$
Since $\pi$ is a covering, each $N_{i}$ is finite and well defined. We also set $N:=\min_{i}N_{i}$. From (41) we know that
$1\leq\max_{x\in B_{1}(p)}\#(\pi^{-1}(x)\cap\tilde{B}_{1}(\tilde{p}))\leq N\leq N_{i},\quad\text{for all }i=1,\ldots,l.$ (42)
Each $\pi^{-1}(A_{i})\cap\tilde{B}_{5}(\tilde{p})$ contains at least $N_{i}$ disjoint isometric copies of $A_{i}$. Hence,
$\int_{\tilde{B}_{5}(\tilde{p})}F(\mathsf{R}_{\tilde{M},k}\vee 0)\,{\rm d}\widetilde{\mathsf{Vol}}\geq\sum_{i=1}^{l}N_{i}\int_{A_{i}}F(\mathsf{R}_{M,k}\vee 0)\,{\rm d}{\mathsf{Vol}}\geq N\int_{B_{1}(p)}F(\mathsf{R}_{M,k}\vee 0)\,{\rm d}\mathsf{Vol}.$
To conclude, it suffices to show that
$N\geq c(K,n)\frac{\mathsf{Vol}(\tilde{B}_{5}(\tilde{p}))}{\mathsf{Vol}(B_{1}(p))}.$ (43)
For every $i\in\{1,\cdots,l\}$, consider the isometric copies of $A_{i}$ contained in $\pi^{-1}(A_{i})$ that intersect $\tilde{B}_{1}(\tilde{p})$. By (42), these are less than $N$ for every $i$. In particular, $\tilde{B}_{1}(\tilde{p})$ is covered by the union of at most $N$ isometric copies of $A_{i}$ as $i$ varies. Hence
$\widetilde{\mathsf{Vol}}(\tilde{B}_{1}(\tilde{p}))\leq\sum_{i=1}^{l}N\,\mathsf{Vol}(A_{i})=N\,\mathsf{Vol}(B_{1}(p)).$
Condition (43) then follows using Bishop-Gromov’s inequality on $\tilde{M}$.
Proof of 2. Let $\bar{p}\in\tilde{B}_{5}(\tilde{p})$ be such that
@@
We claim that it holds
@@
To prove the claim, let $\{x_{i}\}_{i=1}^{l}\subset\tilde{B}_{5}(\tilde{p})$ be such that $\{\tilde{B}_{1}(x_{i})\}_{i=1}^{l}$ is a covering of $\tilde{B}_{5}(\tilde{p})$ while the balls $\{\tilde{B}_{1/5}(x_{i})\}_{i=1}^{l}$ are all disjoint. Since the balls $\{\tilde{B}_{1/5}(x_{i})\}_{i=1}^{l}$ are all disjoint, using the local doubling property under Ricci bounds, it holds
$\widetilde{\mathsf{Vol}}(\tilde{B}_{5}(\tilde{p}))\geq\sum_{i=1}^{l}\widetilde{\mathsf{Vol}}(\tilde{B}_{1/5}(x_{i}))\geq lc(K,n)\widetilde{\mathsf{Vol}}(\tilde{B}_{5}(\tilde{p})).$
Hence, $l\leq c(K,n)$. It follows
@@
which then implies (44) again by the local doubling property.
We now prove item 2. Set $p^{\prime}:=\pi(\bar{p})$. Let $\{A_{1},\cdots,A_{l}\}$ be a partition of $B_{1}(p^{\prime})$ such that each $A_{i}$ is contained in a covering chart of $\pi$. We set
$N_{i}:=\max_{y\in A_{i}}\,\#(\pi^{-1}(y)\cap\tilde{B}_{1}(\bar{p})).$
and $N:=\max_{i}N_{i}$. For every $i$, let $\{\tilde{A}_{i,j}\}_{j=1}^{N_{i}}$ be the isometric copies of $A_{i}$ in $\pi^{-1}(A_{i})$ which have non-empty intersection with $\tilde{B}_{1}(\bar{p})$. Since $\{\tilde{A}_{i,j}\}_{j=1}^{N_{i}}$ covers $\tilde{B}_{1}(\bar{p})$, it holds

### v2

@@ 4 Thin metric spaces
By the previous lemma, a projection of $b_{1}(0)$ on $b_{2}$ lies in $[b_{2}(3R),b_{2}(\mathrm{L}(b_{2})-3R]$. If such projection lies in $[b_{2}(R),b_{2}(3R)]$ or $[b_{2}(\mathrm{L}(b_{2})-3R),b_{2}(\mathrm{L}(b_{2})-R)]$, then $\gamma$ has a point which is $5R$-close to either $b_{2}(3R)$ or $b_{2}(\mathrm{L}(b_{2})-3R)$ (because $X$ is $(R,D)$-thin), otherwise the same conclusion follows by Lemma 4.5. ∎
@@
Let $(X,\mathsf{d})$ be a compact $(R,D)$-thin metric space (see Definition 4.2). Let $l$ be a segment of maximal length in $X$ and assume that $\mathrm{L}(l)\geq 200R$. If $X$ is not contained in a $200R$-neighbourhood of $l$, then there exists $L>0$ and a distance preserving map $\phi:S^{1}_{L}\to X$, such that $X$ is contained in the $D$-neighbourhood of $\phi(S^{1}_{L})$. In particular, $L\geq 50R$.
@@
Consider now a sequence $\phi_{i}\in A$ minimizing the length functional. Each image of $\phi_{i}$ has a projection on $\gamma_{1}$ lying in $I_{1}^{\gamma_{1}}$. Since $X$ is $(R,D)$-thin, the image of each $\phi_{i}$ has a point which is $D$-close to $I_{1}^{\gamma_{1}}$. Hence, all the images of the maps $\phi_{i}$ are contained in a bounded set of $X$. Since the $\phi_{i}$ are all $1$-Lipschitz, by Ascoli-Arzelà theorem, there exists (modulo passing to a subsequence) $\phi_{\infty}$ such that $\phi_{i}\to\phi_{\infty}$ uniformly. By the previous part of the proof, $\phi_{\infty}\in A$. Since the length functional is lower semicontinuous w.r.t. uniform convergence, $\phi_{\infty}$ is an element of minimal length in $A$.
@@
Let $b_{1},b_{2}\in\{\gamma_{1},\gamma_{2},l\}$ with $b_{1}\neq b_{2}$. Let $t^{b_{1}},t^{b_{2}}\in[0,T]$ be the numbers given in the definition of $A$ relative to $\phi_{\infty}$. Modulo reparametrizing $\phi_{\infty}$ and $\gamma_{1}+\gamma_{2}+l$, we can assume that $b_{1}(\mathrm{L}(b_{1}))=b_{2}(0)$ and $t^{b_{1}}\leq t^{b_{2}}$ with $t^{b_{3}}\notin[t^{b_{1}},t^{b_{2}}]$, i.e. $[t^{b_{1}},t^{b_{2}}]$ corresponds to the arc of $S^{1}_{L}$ that connects $t^{b_{1}}$ and $t^{b_{2}}$ without crossing $t^{b_{3}}$.
@@
By definition of $\phi_{\infty}$, the restriction ${\phi_{\infty}}_{|[t^{b_{1}},t^{b_{2}}]}$ is a segment. By Lemmas 4.13 and 4.5, $\pi_{b_{1}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of either $[b_{1}(3R),b_{1}(\bar{t}^{b_{1}})]$ or $[b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$. If the former case happens, ${\phi_{\infty}}([t^{b_{1}},t^{b_{2}}])$ has a point which is $7D$-close to $b_{1}(3R)$, so that it also contains a point $q$ which is $5R$-close to $b_{1}(0)$. By Corollary 4.14, $b_{1}(0)$ is at least $\min\{\mathrm{L}(b_{1}),\mathrm{L}(b_{3})\}+\mathrm{L}(b_{2})/2-20R$ distant from $\phi_{\infty}(t^{b_{2}})$. Hence,
@@
Modulo reparametrizing, we assume $b_{1}(\mathrm{L}(b_{1}))=b_{2}(0)$ and $t^{b_{1}}\leq t^{b_{2}}\leq t^{b_{3}}$ in $[0,T]$. By Step $1$, there exist $t_{1,2}\in[t^{b_{1}},t^{b_{2}}]$ and $t_{2,3}\in[t^{b_{2}},t^{b_{3}}]$, such that $\phi_{\infty}(t_{1,2})$ is $7D$-close to $b_{2}(\mathrm{L}(b_{2})-20R)$ and $\phi_{\infty}(t_{2,3})$ is $7D$-close to $b_{2}(\mathrm{L}(b_{2})+20R)$. Let $\gamma$ be a segment connecting $\phi_{\infty}(t_{1,2})$ and $\phi_{\infty}(t_{2,3})$. By item 2 of Lemma 4.5 and the fact that $X$ is $(R,D)$-thin, either $\pi_{b_{2}}(\gamma)$ intersects $I^{b_{2}}$, or $\gamma$ contains points that are $7D$ close to $b_{2}(R)$ and $b_{2}(\mathrm{L}(b_{2})-R)$. In the latter case, since $\mathrm{L}(b_{2})\geq 200R$, $\gamma$ would not be a segment. Hence, $\pi_{b_{2}}(\gamma)$ intersects $I^{b_{2}}$. It follows that $\gamma+{\phi_{\infty}}_{|S^{1}_{L}\setminus[t_{1,2},t_{2,3}]}$ belongs to $A$. In particular, ${\phi_{\infty}}_{|[t_{1,2},t_{2,3}]}$ is shorter than $\gamma$, so that it is itself a segment.
@@
It follows that $\phi_{\infty}:S^{1}_{L}\to X$ preserves distances. Since $2\pi L>2R$, the fact that $X$ is $(R,D)$-thin implies that $X$ is in the $D$-neighbourhood of $\phi_{\infty}(S^{1}_{L})$. Since $X$ contains a segment of length greater than $200R$, it follows that $L\geq 50R$, concluding the proof. ∎
The next theorem summarizes the results of this section.
@@
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. There exists a $1$-dimensional manifold $I$, possibly with boundary, and a distance preserving map $\phi:I\to X$, such that $X$ is contained in the $200R$-neighbourhood of $\phi(I)$.
@@
A consequence of Theorem 4.16 is that $(R,D)$-thin metric spaces have $1$-Urysohn width bounded from above by a constant $c(R)>0$ depending only on $R$ (see Corollary 4.18). We recall that the notion of Urysohn width was introduced by Gromov in [FillingGromov, widthGromov, LargeGromov].
Definition 4.17.
A metric space $(X,\mathsf{d})$ has $1$-Urysohn width $\leq d$ if there exists a $1$-simplex $Y$, and a continuous map $f:X\to Y$, such that $\mathrm{diam}(f^{-1}(y))\leq d$ for every $y\in Y$.
Corollary 4.18.
Let $(X,\mathsf{d})$ be an $(R,D)$-thin metric space. Then, there exists $c(R)>0$, a $1$-dimensional connected manifold $Y$, and a continuous map $f:X\to Y$, such that $\mathrm{diam}(f^{-1}(y))\leq c(R)$ for every $y\in Y$. In particular, $X$ has $1$-Urysohn width $\leq c(R)$.
Proof.
Let $\phi:I\to X$ be the map given by Theorem 4.16.
Case 1: $I=[0,l)$, with $l\in[0,+\infty]$.
Let $f:X\to\mathbb{R}$ be the distance from $\phi(0)$. Let $x,y\in f^{-1}(t)$ for some $t\geq 0$, and let $t_{x},t_{y}\in[0,l]$ be such that $\phi(t_{x})$ and $\phi(t_{y})$ are projections of $x$ and $y$ on $\phi(I)$. Since $X$ is in the $200R$-neighbourhood of $\phi(I)$, it holds $t_{x},t_{y}\in[0,l]\cap[t-200R,t+200R]$, so that $\mathsf{d}(x,y)\leq 1000R$, which implies $\mathrm{diam}(f^{-1}(t))\leq 1000R$ for every $t\geq 0$.
Case 2: $I=\mathbb{R}$.
Let $B:=B_{1000R}(\phi(0))$ and consider the map $f:X\to\mathbb{R}$ given by
$f(x):=\begin{cases}\mathsf{d}(B,x)&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in }\phi(\mathbb{R}_{+})\\ -\mathsf{d}(B,x)&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in}\phi(\mathbb{R}_{-}).\end{cases}$
First of all, $f$ is well defined, since if $x$ has projections on $\phi(I)$ both in ϕ(+)italic-$\phi(\mathbb{R}_{+})$ and in ϕ()italic-$\phi(\mathbb{R}_{-})$, then it must have a projection in $\phi([-200R,200R])$, so that $x\in B$. We now check that $f$ is continuous. This is trivial at points in $\bar{B}$, so let us assume that $x\notin\bar{B}$ and, without loss of generality, that $x$ has a projection on $\phi(I)$ lying in ϕ(+)italic-$\phi(\mathbb{R}_{+})$. Let $x_{i}\to x$. Since $x\notin\bar{B}$, $x$ has a projection on $\phi(I)$ lying in $\phi([500R,+\infty))$ so that, for $i$ large enough, $x_{i}$ has a projection on $\phi(I)$ lying in ϕ(+)italic-$\phi(\mathbb{R}_{+})$. It follows that $f$ is continuous at $x$, which implies that $f$ is continuous.
Let now $t\in\mathbb{R}$ and let $x,y\in f^{-1}(t)$. If $t=0$, then $\mathsf{d}(x,y)\leq 1000R$. Assume now that $t>0$, and let $t_{x},t_{y}\in\mathbb{R}$ be such that $\phi(t_{x})$ and $\phi(t_{y})$ are projections of $x$ and $y$ on $\phi(I)$. Since $t>0$, it follows that $t_{x},t_{y}>0$, so that $t_{x},t_{y}\in[\mathsf{d}(\phi(0),x)-200R,\mathsf{d}(\phi(0),x)+200R]$, which implies $\mathsf{d}(x,y)\leq 1000R$. The case $t<0$ is analogous, concluding the case $I=\mathbb{R}$.
Case 3: $I=S^{1}_{L}=[0,2\pi L]/\sim$, where $\sim$ identifies the endpoints of $[0,2\pi L]$.
If $L\leq 10^{5}R$, the zero map $f:X\to\mathbb{R}$ defined as $f(x):=0$ gives the desired $1$-Urysohn width estimate. So, we can assume $L\geq 10^{5}R$. Let $B_{1}:=B_{1000R}(\phi(0))$, $B_{2}:=B_{1000R}(\phi(\pi L))$, and consider the map $f:X\to[0,2\mathsf{d}(B_{1},B_{2})]/\sim$, where $\sim$ identifies the endpoints of $[0,2\mathsf{d}(B_{1},B_{2})]$, given by
$f(x):=\begin{cases}\mathsf{d}(B_{1},x)\wedge\mathsf{d}(B_{1},B_{2})&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in }\phi([0,\pi L])\\ (\mathsf{d}(B_{1},B_{2})+\mathsf{d}(B_{2},x))\wedge 2\mathsf{d}(B_{1},B_{2})&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in }\\ &\;\phi([\pi L,2\pi L]).\end{cases}$
The map $f$ is well defined, since if $x$ has projections on $\phi(I)$ both in $\phi([0,\pi L])$ and in $\phi([\pi L,2\pi L])$, then $x\in B_{1}\cup B_{2}$. We now check that $f$ is continuous. This is trivial at points in $B_{1}\cup B_{2}$. So, we consider $x\notin B_{1}\cup B_{2}$ and we assume, without loss of generality, that $x$ has a projection on $\phi(I)$ lying in $\phi([0,\pi L])$. Let $x_{i}\to x$. Since $x\notin B_{1}\cup B_{2}$, $x$ has a projection on $\phi(I)$ lying in $\phi([500R,\pi L-500R])$, so that, for $i$ large enough, $x_{i}$ has a projection on $\phi(I)$ lying in $\phi([0,\pi L])$. It follows that $f$ is continuous at $x$, so that $f$ is continuous.
Let now $t\in[0,2\mathsf{d}(B_{1},B_{2}))$, and let $x\in f^{-1}(t)$. If $t=0$, then $x\in B_{1}$ or $\mathsf{d}(x,B_{2})=\mathsf{d}(B_{1},B_{2})$. It follows that $\mathsf{d}(x,B_{1})\leq 2000R$, showing that $\mathrm{diam}(f^{-1}(0))\leq c(R)$. A similar argument holds for $t=\mathsf{d}(B_{1},B_{2})$. So, let $t\notin\{0,\mathsf{d}(B_{1},B_{2})\}$, and assume without loss of generality that $t\in(0,\mathsf{d}(B_{1},B_{2}))$. Let $x,y\in f^{-1}(t)$, and let $t_{x},t_{y}\in[0,2\pi L)$ be such that $\phi(t_{x})$ and $\phi(t_{y})$ are projections of $x$ and $y$ on $\phi(I)$. Since $t\in(0,\mathsf{d}(B_{1},B_{2}))$, it follows that $t_{x},t_{y}\in[0,\pi L]$. Hence, $t_{x},t_{y}\in[\mathsf{d}(\phi(0),x)-200R,\mathsf{d}(\phi(0),x)+200R]$, so that $\mathsf{d}(x,y)\leq 1000R$. This concludes the proof. ∎
@@ 5 Large scale geometry of almost non-negative Ricci and integrally-positive scalar curvature
5.1 Sufficient curvature conditions for a manifold to be thin
@@
Let $r$ be a segment of length greater than $2R$ in $X$. For every $t\in(R,\mathrm{length}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$, i.e., $(X,\mathsf{d})$ is $(R,D)$-thin (see Definition 4.2).
In particular, there exists a $1$-dimensional connected manifold $I$ (possibly with boundary) and a distance preserving map $\phi:I\to X$ such that $X$ is contained in the $200R$-neighbourhood of $\phi(I)$.
@@
Finally, by Theorem 4.16 , there exists a $1$-dimensional connected manifold $I$ (possibly with boundary) and a distance preserving map $\phi:I\to X$ such that $X$ is contained in the $200R$-neighbourhood of $\phi(I)$. ∎
@@
Let $r$ be a segment of length greater than $2R$ in $X$. For every $t\in(R,\mathrm{length}(r)-R)$ and $x\in\pi_{r}^{-1}(r(t))$, it holds $\mathsf{d}(x,r)<D$, i.e., $(X,\mathsf{d})$ is $(R,D)$-thin (see Definition 4.2).
In particular, there exists a $1$-dimensional connected manifold $I$ (possibly with boundary) and a distance preserving map $\phi:I\to X$ such that $X$ is contained in the $200R$-neighbourhood of $\phi(I)$.
The remaining part of the section is devoted to showing that if a manifold satisfies one of the conditions (34), (35), (36), so do its coverings (modulo a rescaling by dimensional constants). This is the content of Propositions 5.3 and 5.4. The proof of the next proposition follows from [kapovitch2011structurefundamentalgroupsmanifolds, Lemma 1.6]; we report it for completeness of presentation.
@@
Let $n\in\mathbb{N}$ and let $K\in\mathbb{R}$. Let $(M^{n},g,p)$ be a pointed Riemannian manifold with $\mathsf{Ric}_{M}\geq K$. Let $(\tilde{M}^{n},\tilde{g},\tilde{p})$ be the universal covering of $M$, with covering map $\pi:\tilde{M}\to M$ such that $\pi(\tilde{p})=p$. Let $F:M\to\mathbb{R}_{+}$ be a measurable locally bounded function and let $\tilde{F}:=F\circ\pi$.
@@
    $\fint_{\tilde{B}_{3}(\tilde{p})}\tilde{F}\,{\rm d}\widetilde{\mathsf{Vol}}\geq c(K,n)\fint_{B_{1}(p)}F\,{\rm d}\mathsf{Vol}.$
@@
    $\fint_{\tilde{B}_{3}(\tilde{p})}\tilde{F}\,{\rm d}\widetilde{\mathsf{Vol}}\leq c(K,n)\fint_{B_{1}(p^{\prime})}F\,{\rm d}\mathsf{Vol}.$
@@
Consider a measurable function $j:B_{1}(p)\to\tilde{B}_{1}(\tilde{p})$ such that $\pi(j(x))=x$ for every $x\in B_{1}(p)$. Let $T:=j(B_{1}(p))$, so that $\mathrm{diam}(T)\leq 2$ and
@@
Let $S\subset\tilde{M}$ be the union of the sets $g(T)$ for every deck transformation $g:\tilde{M}\to\tilde{M}$ of $\pi:\tilde{M}\to M$ such that $g(T)\cap\tilde{B}_{1}(\tilde{p})\neq\emptyset$. It follows that
@@
Proof of 1. Using that $\widetilde{\mathsf{Vol}}(\tilde{B}_{3}(\tilde{p}))\leq c(K,n)\widetilde{\mathsf{Vol}}(\tilde{B}_{1}(\tilde{p}))$ and (41), it holds
@@
Proof of 2. Arguing as for (41), we find $S^{\prime}\subset\tilde{M}$ such that
@@
Let $\{B_{1}(p_{i})\}_{i\in I}$ be a covering of $B_{3}(p)$ with $I\subset\mathbb{N}$, such that $\{B_{1/5}(p_{i})\}_{i\in I}$ are disjoint. By (42), it holds